# A Dipolar Chiral Spin Liquid on the Breathed Kagome Lattice

### 26 Mar 2026 v1

@@ IV Breathing $\beta$ phase diagram
We can write the two distinct spin-1/2 degrees of freedom as a spin $\vec{s}_{\blacktriangle}$ and a chirality $\vec{\eta}_{\blacktriangle}$ [56]. Whereas $\vec{s}_{\blacktriangle}$ encodes the total spin $\vec{S}_{\text{tot}}$ in this manifold; $\vec{\eta}_{\blacktriangle}$ encodes the spatial spin-correlations within each triangle [Fig. 4(a)]. More specifically, $\hat{\eta}_{\blacktriangle}^{z}$ measures the chirality of the three spins in the triangle, while $\hat{\eta}_{\blacktriangle}^{x/y}$ measures the spatial imbalance of correlations:
@@
$\eta^{y}_{\blacktriangle}$ $=\frac{2}{3}\left[2\vec{S}_{a}\cdot\vec{S}_{b}-\vec{S}_{a}\cdot\vec{S}_{c}-\vec{S}_{b}\cdot\vec{S}_{c}\right],$ (9)
@@
where the right hand side is interpreted before the projection into the appropriate ground state manifold [39].
@@
One might worry that these effective degrees of freedom are only relevant for very large $\beta$ where the system spontaneously breaks translation symmetry. This is not the case: for all $\beta>1$, the ground state’s population in this subspace (per triangle) is $\gtrsim 92\%$ even at the Kagome point $\beta=1.0$ [Fig. 4(b)]. This suggests that these indeed are the relevant low-energy degrees of freedom for describing the underlying phase diagram and phase transitions.
Indeed, the separation of the local Hilbert space between the spin $\vec{s}_{\blacktriangle}$ and the chirality $\vec{\eta}_{\blacktriangle}$ degrees of freedom offers a simple, yet powerful picture for understanding the spontaneous time-reversal symmetry breaking. In this language, time-reversal symmetry breaking is recast as the formation of a long-range ordered Ising state in $\eta^{z}$. At the same time, this identification highlights that the origin of the long-range order in $\chi_{\blacktriangle}$ and $\chi_{\bigtriangledown}$ is very distinct; for $\chi_{\blacktriangle}$ it arises from the low energy chiral modes $\eta^{z}$, while for $\chi_{\bigtriangledown}$ it arises from the correlations within the spin degrees of freedom across different triangles $\chi_{\bigtriangledown}\sim\vec{s}_{\blacktriangle}\cdot(\vec{s}_{\blacktriangle^{\prime}}\times\vec{s}_{\blacktriangle^{\prime\prime}})$.
@@
Having carefully computed the relevant degrees of freedom, we now turn to defining their effective interactions, $H_{\text{eff}}$. We start by recognizing which terms are not allowed by symmetry: the combination of U(1) spin conservation and $\mathbb{Z}_{2}$ spin flip symmetry preclude any single spin $s^{\alpha}_{\blacktriangle}$ terms, while time reversal symmetry also precludes any individual $\eta^{z}_{\blacktriangle}$ [39]. More broadly, because time reversal flips both $\vec{s}$ and $\eta^{z}$, any interactions must have an even number of such terms. By contrast, single- and two-body $\eta^{x/y}_{\blacktriangle}$ terms are not constrained by the onsite symmetries, but rather the lattice symmetries. Interactions between different pairs of sites are related by the rotation and mirror symmetries that define the $C_{3v}$ symmetry group [39]. While these considerations help constrain the different allowed terms, their overall strength remains undetermined. To quantitatively calculate $H_{\text{eff}}$, we need to translate the original spin-spin interactions into triangle-triangle interactions of the effective degrees of freedom.
Figure 4: (a) Schematic of the basis we consider for our $\beta\to\infty$ perturbative effective model. For each triangle composed of 3 spins, we project into the 4 lowest-lying, degenerate states, which can be described by 2 spin-$1/2$ d.o.f.s, $\hat{s}$ and $\hat{\eta}$. $\hat{s}$ encodes the spin rotation properties of this manifold, while $\hat{\eta}$ encodes the chirality and the planar distribution of correlations in the triangle. In our effective model, each pair of spins lives on the vertex of a triangular lattice. (b) Projection of the full wavefunction into this low-energy subspace. Even away from the $\beta\rightarrow\infty$ limit, the overlap of the wavefunction with this subspace remains high ($\gtrsim 92\%$ at the Kagome point $\beta=1$). (c) Using our second-order perturbative effective Hamiltonian $H_{\text{eff}}$, the phase diagram (with respect to $\beta$) of our model closely matches that of the full DMRG calculation for YC8 cylinders with R1 cutoff. This underscores that the $\hat{s}$ and $\eta$ spins are the relevant degrees of freedom that inform the physics across all values of $\beta$.
@@
To zero-th order, $H_{\text{eff}}$ is simply given by the projection $\hat{P}$ of the full system Hamiltonian to the low-energy subspace, $H^{(0)}_{\text{eff}}=\hat{P}H\hat{P}$. We find that $H^{(0)}_{\text{eff}}$ is insufficient to capture the phase diagram of our model. To include higher-order corrections arising from the eliminated states, we perform a Schrieffer–Wolff transformation [39] to obtain an order-by-order perturbative effective Hamiltonian $H_{\text{eff}}=H^{(0)}_{\text{eff}}+H^{(1)}_{\text{eff}}$$\ldots$:
@@
Defining a new two-spin model that includes terms through $H_{\text{eff}}^{(2)}$ and performing iDMRG calculations on this effective model [Fig. 4(c)], we find that the resulting phase diagram exhibits good quantitative agreement with the original system, demonstrating the predictive power of this perturbative approach across all values of $\beta$.
A few open questions remain with respect to the mechanism that stabilizes the CSL. First, at second order, we already obtain an $\eta^{z}\eta^{z}$ Ising term; however, it remains antiferromagnetic across all range of breathings considered, suggesting that the Ising order of $\eta^{z}$ arises from a more complex interplay of different interaction terms.
Second, from symmetry considerations, we expect that our perturbative analysis will generate terms proportional to $\vec{s}_{\blacktriangle}\cdot(\vec{s}_{\blacktriangle^{\prime}}\times\vec{s}_{\blacktriangle^{\prime\prime}})$. While this term is time reversal symmetric, the Ising order in $\eta^{z}$ induces a simple chirality term in the Hamiltonian which is known to stabilize the CSL in frustrated magnets [9]. Indeed, this term is present when we consider our perturbative analysis on three unit-cells, directly highlighting this mechanism. Notably, the iDMRG calculations of Fig. 4 take only into account two-triangle interaction, and thus, do not explicitly include this term in the Hamiltonian, highlighting that the presence of such coupling term is not required for observing the CSL.
V Moving beyond the dipolar XY model
Thus far, we have highlighted how continuous geometric control can be leveraged to drive the system into strongly correlated states such as the chiral spin liquid. A more in-depth understanding of this phase and its stability requires the analysis of the broader phase diagram.
Adding Ising couplings allows us to connect our results to the broader literature studying the CSL in Heisenberg antiferromagnets [8], while the global and/or local magnetic fields provide natural paths to connect the CSL to a trivial paramagnetic phase. This point will become relevant when considering the feasibility of adiabatically preparing the CSL in an AMO platform.
To this end, we focus on three distinct but natural extensions of our original dipolar XY model:
  • (i)
    long-range Ising couplings with strength $\Delta$ between the spins,
  • (ii)
    a global and homogenous magnetic field $h_{x}$ that breaks the system’s U(1)(1) symmetry,
  • (iii)
    a locally varying magnetic field pattern $h_{z}\lambda_{i}$ in the $\hat{z}$ direction, where $\lambda_{i}\in\{-1,1\}$ controls the shape of the pattern.
The most generic Hamiltonian then reads:
$H=H_{XY}+\Delta\sum_{i<j}\frac{S^{z}_{i}S^{z}_{j}}{r^{3}_{ij}}+h_{x}\sum_{i}S^{x}_{i}+h_{z}\sum_{i}\lambda_{i}S^{z}_{i}$ (10)
Role of the Ising coupling $\Delta$
@@ V Moving beyond the dipolar XY model
We begin by investigating the effect of an Ising coupling by taking a variety of cuts in the phase diagram with fixed $\beta$ or $\Delta$. Due to the size of the associated phase space, we restrict our numerics to a fixed geometry (YC8-0) and a fixed interaction cutoff (R1).
@@
The resulting phase diagram [Fig. 5] presents a few intriguing features. First, the CSL phase remains robust across a large area of parameter space, and remains surrounded by either the Kagome Spin Liquid state at small breathings or a translation symmetry breaking state at larger breathings. At the Heisenberg point $\Delta=1$, the CSL is further stabilized, with the system’s critical breathing increasing from $\beta_{c}^{\text{XY}}=2.1(2)$ to $\beta_{c}^{\text{Heis}}=3.0(2)$. Experimentally, this suggests that the stability of the CSL is maintained and perhaps even improved when considering additional Ising van der Waals interactions (relevant in Rydberg tweezer arrays) [57].
@@
Second, the Kagome spin liquid also appears to be quite robust. For all Ising couplings $\Delta$, we continue to observe signatures of the KSL at small breathings; spin-spin correlations remain homogeneous without exhibiting signatures of either long-range order or lattice symmetry breaking (translations or rotations). This observation connects to the expectation in the literature that the easy-axis limit $\Delta\gg 1$ is a Kagome Spin Liquid [23]. Beyond $\Delta\gtrsim 5$, when we no longer observe the CSL state, the phase diagram exhibits a direct transition between the KSL and the translation symmetry broken state. While an intriguing observation, some caveats are important. The KSL is believed to be a gapless critical theory and thus is very sensitive to finite size and interaction range effects. In particular, in finite width cylinders, we expect the underlying gapless mode to become gapped, which might stabilize the phase beyond what occurs in the true thermodynamic behavior. A conclusive understanding of this limit requires a more in-depth and careful study.
@@
Third, when the Ising interaction dominates, the nature of the translation symmetry broken phases is different. In particular, the intervening Glider-VBS phase eventually disappears giving rise to a direct transition between a spin liquid and a VBS state. Moreover, a new AFM-Ising phase dominates the large $\Delta$ and $\beta$ portion of the phase diagram [Fig. 5(inset)].
Role of an external applied field
@@
We now turn to studying the effect of global and local magnetic fields on the phase diagram of our dipolar XY model. These perturbations are particularly well suited for adiabatic preparation protocols: (i) single-site fields are simple to implement and can even be changed dynamically during the experiment, and (ii) for large field strengths, the system’s ground state is an easily preparable paramagnetic state, which serves as an ideal starting point for an adiabatic ramp into the phase of interest [58]. Unfortunately, this approach is far from universal: it requires the two phases (paramagnetic and CSL) to be connected by a continuous, second-order phase transition [59]. Nevertheless, it offers a powerful and robust way of preparing strongly correlated states in a setting where there is no natural bath to cool the system to its ground state.
@@
Figure 6: Phase diagram with respect to a global $\hat{x}$ field of strength $h_{x}$, computed for $\beta=1.5$ on a YC8 cylinder geometry with R1 interaction cutoff. (a) The average magnetization $\expectationvalue{S_{\text{ave}}^{x}}$ as a function of $h_{x}$ displays a series of plateaus consistent with incompressible states at fractional magnetization density. The inset plots display the spatial values for $\langle\sigma_{i}^{x}\rangle$ values for the $\nu=1/3,1/2,\text{ and }2/3$ phases, from left to right. (b) The transition out of the CSL at small $h_{x}$ is first-order, as the observables $\expectationvalue{S_{x}}$ and $\langle\chi_{\blacktriangle}\rangle$ jump discontinuously. This location of this discontinuity depends on the bond dimension considered; upon a finite-size bond dimension scaling analysis, we estimate the true transition to occur at the critical point $h_{x}^{c}\approx 0.0102$, demonstrating the general robustness of the CSL state (inset). (c) At large $h_{x}$ we verify the imcompressible nature of the $\nu=1/2$ and $\nu=2/3$ states by their (approximately) zero susceptibility and finite correlation length. Between them we also observe a reduction in the suceptibility and correlation length, compatible with a more intricate $\nu=7/12$ state.
Global field: $h_{x}$
@@
Local field: $h_{z}$
Broadening the scope of perturbations, we now consider the effect of a locally varying magnetic field pattern, a natural perturbation to consider in the context of emerging AMO platforms [57].
We again focus on breathing $\beta=1.5$, using one particular pattern [highlighted in Fig 1(c)] that is associated with a paramagnetic state with zero total magnetization. Since the Hamiltonian has a $U(1)$ magnetization symmetry, this ensures that by adiabatically ramping down the strength of $h_{z}$, the system can be continuously tuned from the paramagnet to the previously studied zero magnetization CSL state.
@@
Figure 7: Phase diagram with respect to a locally varying magnetic field pattern with strength $h_{z}$. (a) For small $h_{z}$, the system remains in the CSL phase; as $h_{z}$ increases, $\langle\chi_{\blacktriangle}\rangle$ continuously decays, reaching zero at the critical point, $h_{z}^{c}\approx 8.5\times 10^{-3}$. (b) Concomitantly, the correlation length $\xi$ diverges with bond dimension, indicating a second-order transition. Beyond this transition, the system is in a trivial paramagnetic phase. Vertical dashed line denotes the parameter utilized in the paramagnetic state illustrated in Fig. 1d).
VI Probing CSL order in finite-sized clusters
We now turn to analyzing the stability of the CSL ground state in finite-sized clusters, with the goal of understanding the requirements and opportunities to observe its signatures in current AMO experiments.
We begin by searching for evidence of TRS-breaking, a prerequisite of the CSL. To this end, we study the response of the bulk chirality to an applied TRS-breaking global field $H_{\chi}=J_{\chi}\sum_{i}\chi_{\blacktriangle,i}$ [Figs. 8(a,b)]. For clusters with an even number of spins, we observe that the bulk chirality is zero at $J_{\chi}=0$, but exhibits a large, system-size dependent susceptibility. The odd clusters behave drastically differently: for all $J_{\chi}$, the bulk chirality remains nonzero, and exhibits a weak, size-independent response to the $J_{\chi}$ field.
We confirm this interpretation in multiple ways [39]: by directly computing the angular momentum of the cluster and observing its flip when the other chiral ground state is prepared, and by observing the accumulation of magnetization and spin current at the edge of the cluster. We complement this discussion in the Supplementary Material with an in-depth analysis of the symmetry properties of this state [39].
One way to probe whether the system exhibits features of the CSL is to measure its response in a Quantum Hall-like setup. In the context of electrons in the integer or fractional Quantum Hall states, the strength of the magnetic field determines the density of electronic states available, enabling one to tune the electronic density in the sample. The ratio between the changing electronic density $n$ and magnetic flux (per unit cell) is given by the transverse conductivity of the state—this is encoded in Streda’s formula [61]:
$\frac{\Delta n}{\phi/2\pi}=\sigma$ (11)
The same phenomenon occurs in bosonic FQH states of spins (owing to the similar structure of the wavefunction) when the motion of the spins picks up a phase factor upon hopping. Such complex hopping can be engineered in our calculation by including a phase term related to the starting and ending locations of the excitation $S^{+}_{i}S^{-}_{j}\to S^{+}_{i}S^{-}_{j}e^{i\phi_{ij}}$; by choosing the phases appropriately, one can simulate the effect of an homogenous magnetic field of flux $\phi$ [39].
Focusing on the magnetization at the center of the cluster, we observe that the linear relationship of Eq. 11 only occurs for large enough clusters, $N\gtrsim 57$. The slope of this linear increase is also in agreement with our expectations; the extracted spin transversal conductivity $\sigma_{xy}^{N=57}\in[0.39(1),0.69(1)]$ for $\beta\in[1.3,1.5]$ is in agreement with the value of $\sigma_{xy}=1/2$ expected for the CSL state [Fig. 8(d)]. Upon increasing the size of the cluster from 57 to 75, we observe a much more robust response; this highlights that while the observed conductivities are close to the expected value, there are still finite-sized effects that can be ameliorated by increasing the system size.
@@ VI Probing CSL order in finite-sized clusters
Figure 9: Chiral-chiral correlations between selected pairs of triangles for the $N=57$ spin cluster ground state with an applied local magnetic field pattern $h_{z}$. The correlations remain negligible until $h_{z}\lesssim 0.2$, when some pairs start developing correlations. Below $h_{z}\lesssim 7\times 10^{-3}$, all correlations become ferromagnetic, which we use as a proxy for long-range order in a finite system. The red line shows the average chirality over all triangles in the system, which also becomes non-negligible at this field strength. (Inset) Spatial correlations for the cluster at the red dashed line– the spins mostly align with the field. The letters that label the triangles indicate which chiral-chiral correlations we consider in (a).
Experimental Preparation of the CSL cluster
@@
We now turn to considering the preparation requirements of the CSL state by leveraging the second-order phase transition from a staggered local magnetic field [Fig. 7]. To this end, we begin by considering the ground state for the 57 spin cluster as a function of the applied field $h_{z}$ [Fig. 9], monitoring the emergence of long-range order of $\chi_{\blacktriangle}$ as a proxy for the onset of the CSL topological order.

### 15 Apr 2026 v2

@@ II.3 Building an effective spin-chirality model
Figure 4: (a) Schematic of the basis we consider for our $\beta\to\infty$ perturbative effective model. For each triangle composed of 3 spins, we project into the 4 lowest-lying, degenerate states, which can be described by 2 spin-$1/2$ d.o.f.s, $\hat{s}$ and $\hat{\eta}$. $\hat{s}$ encodes the spin rotation properties of this manifold, while $\hat{\eta}$ encodes the chirality and the planar distribution of correlations in the triangle. In our effective model, each pair of spins lives on the vertex of a triangular lattice. (b) Spin correlations for different $\vec{\eta}$ states (with total spin $m_{s}=-1/2$) in the low-energy manifold. (c) Projection of the full wavefunction into this low-energy subspace. Even away from the $\beta\rightarrow\infty$ limit, the overlap of the wavefunction with this subspace remains high ($\gtrsim 92\%$ at the Kagome point $\beta=1$). (d) Using our second-order perturbative effective Hamiltonian $H_{\text{eff}}$, the phase diagram (with respect to $\beta$) of our model closely matches that of the full DMRG calculation for YC8 cylinders with R1 cutoff. This underscores that the $\hat{s}$ and $\eta$ spins are the relevant degrees of freedom that inform the physics across all values of $\beta$.
We can write the two distinct spin-1/2 degrees of freedom as a spin $\vec{s}_{\blacktriangle}$ and a chirality $\vec{\eta}_{\blacktriangle}$ [57]. Whereas $\vec{s}_{\blacktriangle}$ encodes the total spin $\vec{S}_{\text{tot}}$ in this manifold; $\vec{\eta}_{\blacktriangle}$ encodes the spatial spin-correlations within each triangle [Fig. 4(a)]. More specifically, $\hat{\eta}_{\blacktriangle}^{z}$ measures the chirality of the three spins in the triangle, while $\hat{\eta}_{\blacktriangle}^{x/y}$ measures the spatial imbalance of correlations [Fig. 4(b)]:
@@
$\eta^{y}_{\blacktriangle}$ $=\frac{2}{3}\left[\vec{S}_{a}\cdot\vec{S}_{c}+\vec{S}_{b}\cdot\vec{S}_{c}-2\vec{S}_{a}\cdot\vec{S}_{b}\right],$ (8)
@@
where the right-hand side is interpreted before the projection into the appropriate ground state manifold [41].
@@
One might worry that these effective degrees of freedom are only relevant for very large $\beta$ where the system spontaneously breaks translation symmetry. This is not the case: for all $\beta>1$, the ground state’s population in this subspace (per triangle) is $\gtrsim 92\%$ even at the Kagome point $\beta=1.0$ [Fig. 4(b)]. This suggests that these are the relevant low-energy degrees of freedom for describing the underlying phase diagram and phase transitions.
Reframing of the original spin-1/2 system into $\vec{s}$ and $\vec{\eta}$ degrees of freedom offers two key advantages: (i) it provides a more effective basis that improves the convergence of iDMRG calculations, and (ii) it offers a new language to probe the mechanism that stabilizes the CSL.
@@
To directly compute the correlated ground state in the $\vec{s}$ and $\vec{\eta}$ basis, we must first calculate their effective interactions, $H_{\text{eff}}$. We start by recognizing which terms are not allowed by symmetry: the combination of U(1) spin conservation and $\mathbb{Z}_{2}$ spin flip symmetry preclude any single spin $s^{\alpha}_{\blacktriangle}$ terms, while time reversal symmetry also precludes any individual $\eta^{z}_{\blacktriangle}$ [41]. More broadly, because time reversal flips both $\vec{s}$ and $\eta^{z}$, any interactions must have an even number of such terms. By contrast, single- and two-body $\eta^{x/y}_{\blacktriangle}$ terms are not constrained by the onsite symmetries, but rather the lattice symmetries. Interactions between different pairs of sites are related by the rotation and mirror symmetries that define the $C_{3v}$ symmetry group [41].
@@
To zero-th order, $H_{\text{eff}}$ is simply given by the projection $\hat{P}$ of the full system Hamiltonian to the low-energy subspace, $H^{(0)}_{\text{eff}}=\hat{P}H\hat{P}$. Crucially, we find that $H^{(0)}_{\text{eff}}$ is insufficient to capture the phase diagram of our model as it contains only a small subset of all possible terms. Higher-order corrections arise from virtual processes that are mediated by the eliminated states; using a Schrieffer–Wolff transformation [41] we obtain an order-by-order perturbative effective Hamiltonian $H_{\text{eff}}=H^{(0)}_{\text{eff}}+H^{(1)}_{\text{eff}}$$\ldots$:
@@
Defining a new two-spin model that includes terms through $H_{\text{eff}}^{(2)}$ and performing iDMRG calculations on this effective model [Fig. 4(d)], we find that the resulting phase diagram exhibits good quantitative agreement with the original system, demonstrating the predictive power of this perturbative approach across all values of $\beta$. Notably, this approach allows us to capture the first-order phase transition into the Glider VBS with smaller bond dimension; local updates on the iDMRG optimization affect tensors corresponding to six spins $\vec{S}$, helping to establish the Glider VBS pattern of correlations more effectively.
More broadly, the separation of the local Hilbert space between the spin $\vec{s}_{\blacktriangle}$ and the chirality $\vec{\eta}_{\blacktriangle}$ degrees of freedom offers a simple, yet powerful picture for understanding the spontaneous time-reversal symmetry breaking. Three observations are important. First, time-reversal symmetry breaking can be simply recast as the formation of a long-range ordered Ising state in $\eta^{z}$. Second, the nature of the long-range order in $\chi_{\blacktriangle}$ and $\chi_{\bigtriangledown}$ is very distinct. For $\chi_{\blacktriangle}$, it arises from ordering of the local chiral $\eta^{z}$ d.o.f.s, while for $\chi_{\bigtriangledown}$, it arises from ordering of complex three-body spin correlations across different triangles $\chi_{\bigtriangledown}\sim\vec{s}_{\blacktriangle}\cdot(\vec{s}_{\blacktriangle^{\prime}}\times\vec{s}_{\blacktriangle^{\prime\prime}})$. Third, from symmetry, $H_{\text{eff}}$ can contain a term proportional to $\vec{s}_{\blacktriangle}\cdot(\vec{s}_{\blacktriangle^{\prime}}\times\vec{s}_{\blacktriangle^{\prime\prime}})$. While this term is time reversal symmetric, when $\eta^{z}$ become ordered, this term induces a simple spin chirality term in the Hamiltonian—this term is known to stabilize the CSL in frustrated magnets [9]. Indeed, by extending our Schrieffer-Wolff approach to three unit-cells, we observe this term in $H_{\text{eff}}$.
However, the mechanism driving the CSL state appears to be more robust than expected. While we observe the Ising order of $\eta^{z}$, we find that the effective Hamiltonian does not ferromagnetically couple $\eta^{z}\eta^{z}$, but rather anti-ferromagnetically couples them. This remains true even when considering terms mediated by spin-spin correlations (e.g. $(s^{x}_{\blacktriangle}s^{x}_{\blacktriangle^{\prime}}+s^{y}_{\blacktriangle}s^{y}_{\blacktriangle^{\prime}})\eta^{z}_{\blacktriangle}\eta^{z}_{\blacktriangle^{\prime}}$). Even more surprisingly, while the $\vec{s}_{\blacktriangle}\cdot(\vec{s}_{\blacktriangle^{\prime}}\times\vec{s}_{\blacktriangle^{\prime\prime}})$ term offers a natural mechanism for the CSL state, we find that the CSL remains robust even when this term is not included. Indeed, the iDMRG calculations of Fig. 4 only take into account two-triangle interactions, and thus, do not include this three-triangle interaction. These two facts highlight how the CSL state is much more robust than expected and driven by a more nuanced mechanism. Overall, we hope that the $\vec{s}$ and $\vec{\eta}$ degrees of freedom provide a novel perspective and help to better study its origin.
III Experimental Considerations
In this section we explore how the ingredients of modern AMO platforms—namely the ability to design the system’s interaction, geometry, and apply generic fields—can be leveraged to better stabilize, prepare and probe the CSL.
III.1 Stability
So far, we have only studied the stability of the CSL state in an infinite geometry, using dipolar interactions. To connect with a suitable experimental realization, we need to consider the stability of the phase in finite-sized clusters. On the one hand, such investigation provides a direct measure of the extent of the experiment required to see this topological state [Fig. 5]. On the other hand, it offers the opportunity to directly explore the edge mode via spatial correlations rather than through the entanglement spectrum of in the infinite cylinder geometry.
While finite system sizes can hurt, engineering different interactions using Floquet engineering offers a path towards further stabilizing the CSL state [58, 59, 60]. To this end, we explore the broader breathing phase diagram upon the inclusion of a long-range dipole Ising term [Fig. 6].
Finite Clusters
Much like in the infinite geometry state, we begin by studying the stability of the CSL state in finite clusters by searching for evidence of TRS-breaking. To this end, we study the response of the bulk chirality to an applied TRS-breaking global field $H_{\chi}=J_{\chi}\sum_{i}\chi_{\blacktriangle,i}$ [Figs. 5(a,b)]. For clusters with an even number of spins, we observe that the bulk chirality is zero at $J_{\chi}=0$, but exhibits a large, system-size dependent susceptibility. The odd clusters behave drastically differently: for all $J_{\chi}$, the bulk chirality remains nonzero, and exhibits a weak, size-independent response to the $J_{\chi}$ field.
We confirm this interpretation in multiple ways [41]: by directly computing the angular momentum of the cluster and observing its flip when the other chiral ground state is prepared, and by observing the accumulation of magnetization and spin current at the edge of the cluster.
As we saw in Section II.1, the TRS-breaking cannot unambiguously characterize the CSL; one must also probe nonlocal quantities. One way to probe whether the system exhibits features of the CSL is to measure its conductivity, or response in a Quantum Hall-like setup. The ratio between the changing electronic density $n$ and magnetic flux (per unit cell) is given by the transverse conductivity of the state—this is encoded in Streda’s formula [61]:
$\frac{\Delta n}{\phi/2\pi}=\sigma$ (9)
The magnetic flux can be engineered by including a complex phase in the hopping term related to the starting and ending locations of the excitation $S^{+}_{i}S^{-}_{j}\to S^{+}_{i}S^{-}_{j}e^{i\phi_{ij}}$; by choosing the phases appropriately, one can simulate the effect of an homogenous magnetic field of flux $\phi$ [41].
Focusing on the magnetization at the center of the cluster, we observe that the linear relationship of Eq. 9 only occurs for large enough clusters, $N\gtrsim 57$. The slope of this linear increase is also in agreement with our expectations; the extracted spin transverse conductivity $\sigma_{xy}^{N=57}\in[0.39(1),0.69(1)]$ for $\beta\in[1.3,1.5]$ is in agreement with the value of $\sigma_{xy}=1/2$ expected for the CSL state [Fig. 5(d)]. Upon increasing the size of the cluster from 57 to 75, we observe a much more robust response; this highlights that while the observed conductivities are close to the expected value, there are still finite-sized effects that can be ameliorated by increasing the system size.
Ising coupling $\Delta$
Ising interactions arise in dipolar XY tweezer arrays in two ways: (i) the natural van der Waals $1/r^{6}$ coupling and (ii) Floquet engineering. This expands the landscape of possible Hamiltonians one can consider by carefully designing the absolute length scale of the Kagome lattice, or periodically modulating the dipolar system.
@@ III.1 Stability
We now investigate how this term affects the stability of the CSL by considering a variety of cuts in the phase diagram with fixed $\beta$ or $\Delta$. Due to the size of the associated phase space, we restrict our numerics to a fixed geometry (YC8-0) and a fixed interaction cutoff (R1).
@@
The resulting phase diagram [Fig. 6] presents a few intriguing features. First, the CSL phase remains robust across a large area of parameter space. At the Heisenberg point $\Delta=1$, the CSL is further stabilized, with the critical breathing increasing from $\beta_{c}^{\text{XY}}=2.1(2)$ to $\beta_{c}^{\text{Heis}}=3.0(2)$. Experimentally, this suggests that the stability of the CSL is maintained and perhaps even improved when considering additional Ising van der Waals interactions [62].
@@
Second, the Kagome spin liquid also appears to be quite robust. For all Ising couplings $\Delta$, we continue to observe signatures of the KSL at small breathings; spin-spin correlations remain homogeneous without exhibiting signatures of either long-range order or lattice symmetry breaking (translations or rotations). This observation connects to the expectation in the literature that the easy-axis limit $\Delta\gg 1$ is a Kagome Spin Liquid [24]. Beyond $\Delta\gtrsim 5$, the phase diagram exhibits a direct transition between the KSL and the translation symmetry broken state. While an intriguing observation, some caveats are important. The KSL is believed to be a gapless critical theory and thus is very sensitive to finite size and interaction range effects. In particular, in finite width cylinders, we expect the underlying gapless mode to become gapped, which might stabilize the phase beyond what occurs in the true thermodynamic behavior. A conclusive understanding of this limit requires a more in-depth and careful study.
@@
Third, when the Ising interaction dominates, the nature of the translation symmetry broken phases is different. At large $\beta$ and $\Delta$, we observe a new antiferromagnetic phase (AFM-Ising) [Fig. 6(inset)].
III.2 Preparation
Having established the stability of the CSL state, we now address its preparation in a quantum simulator. Because the CSL is a topological state, preparation using unitary dynamics must scale with the system size [63]. However, unlike other topological states previously studied in these devices, the lack of a zero-correlation-length reference state for this phase prevents the development of digital approaches [64].
To this end, we consider the ingredients for the adiabatic preparation of the CSL state. Such approach offers guarantees for the preparation of the ground state supported by the long-range interactions considered in our work, provided we can: (i) prepare a Hamiltonian and its ground state with high fidelity, and (ii) continuously deform the Hamiltonian to the CSL Hamiltonian [65].
Crucially, these conditions are not met by our breathing phase diagram. Preparing from the Kagome Spin liquid state requires the preparation of this complex state—an outstanding challenge in its own right [66, 67]. At the same time, preparing from the valence bond solid is blocked by the presence of a discontinuous transition which precludes a smooth deformation of the ground state across the phase transition.
As a result, we focus on the effect of global and local fields on the breathing phase diagram of the form:
$H_{\text{global}}=h_{x}\sum_{i}S^{x}_{i}\quad\text{and}\quad H_{\text{local}}=h_{z}\sum_{i}\lambda_{i}S^{z}_{i},$ (10)
where $\lambda_{i}=\pm 1$ determines the pattern of the local perturbation.
@@ III.2 Preparation
These perturbations are particularly well suited for adiabatic preparation protocols. Single-site fields are simple to implement and can even be changed dynamically during the experiment, and at large field strengths, the system’s ground state is an easily preparable paramagnetic state, which serves as an ideal starting point for an adiabatic ramp into the phase of interest [17].
@@
Figure 7: Phase diagram with respect to a global $\hat{x}$ field of strength $h_{x}$, computed for $\beta=1.5$ on a YC8 cylinder geometry with R1 interaction cutoff. (a) The average magnetization $\expectationvalue{S_{\text{ave}}^{x}}$ as a function of $h_{x}$ displays a series of plateaus consistent with incompressible states at fractional magnetization density. The inset plots display the spatial values for $\langle\sigma_{i}^{x}\rangle$ values for the $\nu=1/3,1/2,\text{ and }2/3$ phases, from left to right. (b) The transition out of the CSL at small $h_{x}$ is first-order, as the observables $\expectationvalue{S_{x}}$ and $\langle\chi_{\blacktriangle}\rangle$ jump discontinuously. This location of this discontinuity depends on the bond dimension considered; upon a finite-size bond dimension scaling analysis, we estimate the true transition to occur at the critical point $h_{x}^{c}\approx 0.0102$, demonstrating the general robustness of the CSL state (inset). (c) At large $h_{x}$ we verify the rigid nature of the $\nu=1/2$ and $\nu=2/3$ states by their (approximately) zero susceptibility and finite correlation length. Between them we also observe a reduction in the susceptibility and correlation length, compatible with a more intricate $\nu=7/12$ state.
Global field $h_{x}$
@@
Local field $h_{z}$
In general, many different patterns can be achieved by changing the values of $\lambda_{i}$ in Eq. 10. Here, we focus on a specific one—corresponding to the paramagnetic state in Fig 8(b)—which minimizes the total magnetization in each triangle, and globally favors the state of zero magnetization. Since the Hamiltonian has a $U(1)$ spin rotation symmetry, magnetization is conserved as we adiabatically ramp down the strength of $h_{z}$, connecting the initial paramagnetic state to the previously studied zero magnetization CSL state.
@@
Figure 8: Phase diagram with respect to a locally varying magnetic field pattern with strength $h_{z}$. (a) For small $h_{z}$, the system remains in the CSL phase; as $h_{z}$ increases, $\langle\chi_{\blacktriangle}\rangle$ continuously decays, reaching zero at the critical point, $h_{z}^{c}\approx 8.5\times 10^{-3}$. (b) Concomitantly, the correlation length $\xi$ diverges with bond dimension, indicating a second-order transition. Beyond this transition, the system is in a trivial paramagnetic phase. Vertical dashed line denotes the parameter utilized in the paramagnetic state illustrated in Fig. 1(d). (c) Chiral-chiral correlations between selected pairs of triangles in an $N=57$ spin cluster. The correlations remain negligible until $h_{z}\lesssim 0.2$, when some pairs start developing correlations. Below $h_{z}\lesssim 7\times 10^{-3}$, all correlations become ferromagnetic, which we use as a proxy for long-range order in a finite system. The red line shows the average chirality over all triangles in the system, which also becomes non-negligible at this field strength. (d) Spatial correlations for the cluster at the red dashed line– the spins mostly align with the field. The letters that label the triangles indicate which chiral-chiral correlations we consider in (a).
Preparing Finite Clusters
@@
We now turn to considering the preparation requirements of the CSL state by leveraging the second-order phase transition from a staggered local magnetic field [Fig. 8]. To this end, we begin by considering the ground state for the 57 spin cluster as a function of the applied field $h_{z}$, monitoring the emergence of long-range order of $\chi_{\blacktriangle}$ as a proxy for the onset of the CSL topological order.